Please help me solve this question:
"The ages of three sisters are three consecutive even integers. The sum of the youngest and the oldest sisters' ages is twenty less than 4 times the middle sisters age. What is the age of the middle sister? Only an algebraic solution will be accepted."

Answer by richwmiller(17219)

a
a+2
a+4
a+a+4=4*(a+2)-20
2a+4=4a+8-20
2a+4=4a-12
2a=16
a=8
a+2=10
a+4=12